Problem 3 (10 pts). Let f(x)-δ(z + a) + δ(z-a); Ict r(z) and h(z) be functions de- scribed in Fig. 2 below. As discussed in class, one can show thatof(u)r u)du Assume that f(x) and h(x) are known but r(x) has been lost. Recover r(x) f(x) r(x) h(x) co 1f -bl b -a FIG. 2: This refers to Problem 3.
Answer #1

r(x) can be obtained by the scaled convolution of f(x) and h(x).

r(x)=rac{1}{2} int_{-infty}^{+infty} h(u)f(x-u)du

graph of f(x-u)

f(x-u) -a-x a-x

for x<-b

The overlapping graph of f(x-u) with h(u) is

from the graph, it is evident that the value of the integral

r(z) =- [h(-a-x) + h(a-r)] = 210 + 0] = 0

for this case (x<-b) is 0.

_______________________________________________________________________________

case ii : -b<x<b

-a-x a-x a -a

The value of integral is

r(z) =-

________________________________________________________________________-

for the case when x>b

-a-x a-x -a

r(z) =- [h(-a-x) + h(a-r)] = 210 + 0] = 0

___________________________________________________________________

Hence ,

from the three cses discussed above

r(x)=0 for x <-b r(z) = 1 for-b<x<b r(x)=0 for b<x

Hence we recover r(x)

r(x) -b